40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 4 00:10:43 2020, gautam, Update, General, Vent activities today IMG_0184.jpgIMG_0185.jpgIMG_0187.jpgIMG_0207.jpegIMG_0210.jpg
    Reply  Thu Nov 5 22:27:34 2020, gautam, Update, General, Vent activities today 
       Reply  Fri Nov 6 11:36:37 2020, gautam, Update, General, ETMY suspension eigenmodes ETMY_pkFitNaive.pdfETMY.tar.bz2
          Reply  Tue Nov 10 15:13:41 2020, rana, Update, General, ETMY suspension eigenmodes 
             Reply  Thu Nov 12 14:26:35 2020, gautam, Update, General, ETMY suspension eigenmodes ETMY.tar.bz2
       Reply  Sun Nov 8 20:22:35 2020, gautam, Update, General, Vent activities today IFOalignment.png
          Reply  Sun Nov 8 23:24:29 2020, Koji, Update, General, Vent activities today 
          Reply  Mon Nov 9 17:40:17 2020, gautam, Update, General, Pumpdown IMG_8937.jpgpumpdown.png
             Reply  Tue Nov 10 11:31:13 2020, Koji, Update, General, Pumpdown Screen_Shot_2020-11-10_at_11.30.21.png
Message ID: 15660     Entry time: Thu Nov 5 22:27:34 2020     In reply to: 15658     Reply to this: 15661   15664
Author: gautam 
Type: Update 
Category: General 
Subject: Vent activities today 

[koji, gautam]

  1. Tweaked last steering mirror before PSL beam is launched into vacuum to get the IMC resonant for the TEM00 mode. Less than 1/8 turn in each Pitch and Yaw was required, and we recovered MC2T ~ 1300 cts (c.f. the expected ~1500 cts when the cavity is well aligned, but we didn't touch the IMC mirrors).
  2. ETMY magnet re-gluing
    • Regluing was successful. Pickle picker came off smoothly. 
    • We performed the razor blade test on all 6 magnets, the integrity of the joints seems uncompromised.
  3. ETMY cleaning
    • Applied F.C. to HR face, HR edge (to remove some residual stains) and AR surface (taking care to go around the magnets).
    • A fresh bottle each of spectroscopic grade acetone and isopropanol was dispensed into clean beakers for the work.
    • Curiously, in the lighting conditions of the cleanroom, both HR and AR faces looked surprisingly clean, even when illuminated with the green flashlight.
    • However, when we took it to the east end and looked at it again with the green flashlight, just before putting it into the chamber, we saw all sorts of stains and markings on the HR side.
    • The central aperture looks fine - the contamination did not happen during the transport so we feel confident not to futz around more with this and since we cleaned it extensively in the cleanroom, we opted not to do another round of cleaning in the chamber.
  4. ETMY re-suspension (on the existing wire loop)
    • The back OSEM-holding plates were removed for access reasons.
    • The optic was rested onto the EQ stops. 
    • Patient nudging of the wire around the optic allowed us to eventually get the wire into the V-grooves in the standoffs.
    • The wire was not damaged! (at least, to eye, free-swinging test is going on now, we will see what the eigenfrequencies and Qs are)...
    • We confirmed that the wire is seated in the v-groove, and took some close up photos.
  5. ETMY pitch balancing
    • A HeNe was brought into the cleanroom and mounted at 5.5" beam height.
    • Level-ness of the beam (i.e. parallel to the optic table) was done coarsely with a spirit level placed on the HeNe holder, and more fine adjustment was done by checking the beam height at the launch point, and again ~2 meters away.
    • All EQ stops were backed off to ensure the optic was really "free-swinging" - no OSEMs either at this point, since all this was done in the cleanroom.
    • The return beam hit the HeNe's output aperture, so we were satisfied that the pitch balancing was good to ~1mrad.
  6. ETMY transport back to EY
    • Since we want to maintain some tension in the wires, the bottom EQ stops were not raised as much as they were when the tower was transported from EY to the cleanroom.
    • As a result, we felt it'd be better for a human (Koji) to carry the cage (rather than cart it, the idea being that there would be fewer vertical impulses due to bumps on the ground etc).
    • The transport went smooth.
  7. EY chamber work
    • Optic was placed back in its original position (marked by some L-clamps).
    • In preparation for alignment work, I undid all the CDS changes I made to facilitate the temporary 3-magnet actuation. This necessitated a reboot of c1auxey VME crate.
    • OSEMs were inserted - best effort to half light as usual, and we tried to replicate the rotation in the mounts as closely as possible to what was the case before (using a photo from the last vent as reference).
    • Coarse alignment was done by making the Oplev return beam hit the QPD.
    • Better alignment of the Y-arm cavity axis was done using the green beam. 
    • IMC was locked - the beam alignment target was used to ensure the IR beam was hitting ETMY at the correct height (by moving TT2), and then the ETMY angular alignment was improved to make the return beam go back roughly collinearly.
  8. Y-arm POY locking
    • Satisfied with the alignment, we returned to the control room (light doors back on EY chamber) - cavity alignment was tweaked further to make the TRY resonances as high as possible.
    • I could lock the Y-arm length to the PSL frequency using POY as a sensor - TRY was ~0.04-0.05 (since IMCT is 1/10th its usual value, we expect it to be 0.1, so this is not bad).
  9. OMC chamber work
    • Tied down the balance weights that were previously placed after the OMC and peripheral optics were removed.
    • Checked the table leveling.
    • Checked that the AS beam is reasonably well centerd on OM5 and OM6. Took some photos.
    • Checked that the IMC could be locked after this work - it could.

So all the primary vent objectives have been achieved 🙌 . The light doors are on the chamber right now. I'm measuring the free-swinging spectra of ETMY overnight. Barring any catastrophic failures and provided all required personnel are available, we will do the final pre-close-up checks, put the heavy doors back on, and pump down starting 10 am Monday, 9 Nov 2020. Some photos here.

ELOG V3.1.3-